There are four numbers such that first three of them from an A.P. and the last three form a G.P. The sum of the first and third number is 2 and that of second and fourth is 26. What are these numbers ?


1
Expert's answer
2021-10-26T15:23:15-0400
  1. let the 1st term be (a-d), a, (a+d), where a is the 1st term and d be the common difference a-d+a+d=2, then 2a=2 "\therefore" a=2. Given a+ar2=26 ,factor out a "\to" a(1+r2)=26 but a=1 "\therefore" r2=25"\to" r=5 or -5 2nd term=a,3rd term=ar=5 or -5,4th term =ar2=25. d=a3-a2=5-1=4 or -5-1=-6, "\therefore" we know that a1=a2-d=1-4=-3 or1-(-6)=7 "\therefore" there are two sets of numbers {-3,1,5,25} and {7,1,-5,25}